### v1

@@ 5 Conclusions and Future Work
This article introduced Rhino-Cyt as an EUD system for AI-assisted medical diagnostics, proposing explanation-driven intervention as a novel approach to AI customization. Traditional AI customization methods often limit the role of domain experts to passive reviewers. Rhino-Cyt, by contrast, enables experts to directly manipulate AI-generated outputs and their explanations, offering a way to refine AI behavior without requiring any programming expertise. Rhino-Cyt broadens the scope of AI customization and fosters trust calibration, ensuring that model outputs remain aligned with experts’ reasoning. Through its positioning within existing EUD taxonomies (2), this study highlighted Rhino-Cyt’s hybrid nature, sitting at the intersection of XAI and user-driven model customization. The three-step interaction workflow—classification overrides, explanation editing, and impact visualization—aims to allow domain experts to progressively adapt AI systems while maintaining meaningful control over decision-making processes. This work contributes to ongoing research in human-centered AI and AI-assisted decision support, demonstrating that intervention-based, explainability-driven EUD systems have the potential to bridge the gap between AI automation and expert oversight. By shifting AI-assisted diagnostics toward a model where human expertise plays an active role in shaping AI behavior, Rhino-Cyt paves the way for more transparent, trustworthy, and adaptable AI systems. The open questions regarding Rhino-Cyt and its implementation of explanation-driven interventions concern their usability, effectiveness, and broader applicability. The first step will be conducting a study with users (i.e., rhinocytologists) to assess its usability and determine its strengths and weaknesses to improve in the next development iteration. One key direction for future research is expanding its capabilities beyond single-user interactions. Medical diagnostics often rely on expert consensus, and extending Rhino-Cyt to support multi-user collaboration could enable a more robust AI adaptation process. Shared knowledge bases and federated learning could further enhance this by allowing explanation refinements to inform AI behavior across multiple institutions, ensuring that AI systems continuously learn from diverse expert insights. Another critical area of investigation is the impact of editable explanations on AI performance. Future studies should assess how user interventions influence model accuracy, interpretability, and cognitive load. A structured evaluation with domain experts could provide insights into whether explanation-driven interventions foster greater trust and understanding compared to traditional rule-based AI customization. Beyond the medical domain, the principles behind Rhino-Cyt could be applied to other fields, such as legal and financial AI, where professionals must interpret and refine AI-generated reasoning. As AI continues to integrate into high-stakes decision-making, ensuring that users can meaningfully shape its outputs will be essential for fostering reliable, human-centered AI systems. {credits}
@@ 5.0.1 Acknowledgements
This research is partially supported by the co-funding of the European Union - Next Generation EU: NRRP Initiative, Mission 4, Component 2, Investment 1.3 – Partnerships extended to universities, research centers, companies and research D.D. MUR n. 341 del 15.03.2022 – Next Generation EU (PE0000013 – “Future Artificial Intelligence Research – FAIR” - CUP: H97G22000210007). This research is also partially supported by the Italian Ministry of University and Research (MUR) and by the European Union - NextGenerationEU, Mission 4, Component 2, Investment 1.1, under grant PRIN 2022 PNRR “PROTECT: imPROving ciTizEn inClusiveness Through Conversational AI” (Grant P2022JJPBY) - CUP: H53D23008150001. The research of Andrea Esposito is funded by a Ph.D. fellowship within the framework of the Italian “D.M. n. 352, April 9, 2022”- under the National Recovery and Resilience Plan, Mission 4, Component 2, Investment 3.3 – Ph.D. Project “Human-Centred Artificial Intelligence (HCAI) techniques for supporting end users interacting with AI systems,” co-supported by “Eusoft S.r.l.” (CUP H91I22000410007).

### v2

@@ 5 Conclusions and Future Work
This article proposed explanation-driven interventions as an EUD tool for black-box AI systems, illustrated through their adoption in Rhino-Cyt, a system aiding rhinocytologists in cell counting tasks. Unlike traditional AI customization methods that relegate domain experts to passive reviewers, explanation-driven interventions empowers them to directly manipulate AI-generated outputs and explanations, without any programming expertise. This work contributes to ongoing research in human-centered AI and AI-assisted decision supports EUD systems have the potential to bridge the gap between AI automation and expert oversight. Future work includes evaluating the usability and effectiveness of Rhino-Cyt through user studies, identifying improvement areas, and extending support to multi-user collaboration. Better methods for tracking intervention impact and mitigating user errors are also needed. Another key direction is assessing how editable explanations affect model performance, accuracy, and user cognitive load. Structured evaluations could clarify whether this approach enhances trust and understanding more than traditional rule-based customization. Finally, explanation-driven interventions hold promise beyond medicine, with potential in domains like law and finance where professionals must interpret and refine AI reasoning.
{credits}
@@ 5.0.1 Acknowledgements
This research is partially supported by: (i) the co-funding of the European Union - Next Generation EU: NRRP Initiative, Mission 4, Component 2, Investment 1.3 – Partnerships extended to universities, research centers, companies and research D.D. MUR n. 341 del 15.03.2022 – Next Generation EU (PE0000013 – “Future Artificial Intelligence Research – FAIR” - CUP: H97G22000210007); (ii) the Italian Ministry of University and Research (MUR) and by the European Union - NextGenerationEU, Mission 4, Component 2, Investment 1.1, under grant PRIN 2022 PNRR “PROTECT: imPROving ciTizEn inClusiveness Through Conversational AI” (Grant P2022JJPBY) - CUP: H53D23008150001. The research of Andrea Esposito is funded by a Ph.D. fellowship within the framework of the Italian “D.M. n. 352, April 9, 2022”- under the National Recovery and Resilience Plan, Mission 4, Component 2, Investment 3.3 – Ph.D. Project “Human-Centred Artificial Intelligence (HCAI) techniques for supporting end users interacting with AI systems,” co-supported by “Eusoft S.r.l.” (CUP H91I22000410007). The research of Francesco Greco is funded by a Ph.D. fellowship within the framework of the Italian “D.M. n. 352, April 9, 2022”- under the National Recovery and Resilience Plan, Mission 4, Component 2, Investment 3.3 – Ph.D. Project “Investigating XAI techniques to help user defend from phishing attacks”, co-supported by “Auriga S.p.A.” (CUP H91I22000410007).